OpenStudy (anonymous):

A hexagon has area 90 in^2 and shortest side length 5 in. A similar hexagon has shortest side of length 4 in. What is the area of the similar hexagon?

OpenStudy (p0sitr0n):

5/4 = k A1/A2 = k^2 = 25/16 A2 = 57.6